OpenStudy (anonymous):

By what number must you multiply each side of the equation 2/3x=10 to produce the equivalent equation x = 15?

OpenStudy (anonymous):

is it \[\frac23 x=10?\]

OpenStudy (anonymous):

in order to get rid of the fraction 2/3 you must multiply both sides by the reciprocal which is 3/2. sooo.... (3/2)(2/3)x=(3/2)10 the left side will equal 1x and then the right side is 30/2, which is 15 so x=15